The binary operation * is defined by x*y=xy-y-x for all real values of x and y. If x*3=2*x, find x.
Answer by josgarithmetic(39627)

We seem to need to assume, "-" means minus, and ab means a multiplied by b.


x*3=2*x is given and want to solve for x.

Left side, (x)(3)-3-x.
Right side, 2x-x-2.
Left side must equal right side, ...
(x)(3)-3-x=2x-x-2
Combining steps,
3x-x-3=x-2
2x-3=x-2
2x-x-3=x-x-2, just using the additive inverse of x
x-3=-2
x=-2+3, used additive inverse of -3
ANSWER: x=1